hey,
could you please solve the following two problem for me, and thanks,
1. What number and its square differ by 20?
2. The hypotenuse of a right triangle is 29 cm. If the other two sides differ
by 1 cm, what are their lengths?
i will be waiting for the solutions,
and thank you again.

Answer by checkley77(12844)

x^2-x=20
x^2-x-20=0
(x-5)(x+4)=0
x-5=0
x=5 answer.
Proof:
5^2-5=20
25-5=20
-----------------------
x^2+(x+1)^2=29^2
x^2+x^2+2x+1=841
2x^2+2x+1-641=0
2x^2+2x-840=0
2(x^2+x-420)=0
2(x+21)(x-20)=0
x-20=0
x=20 answer.
20+1=21 answer.
Proof:
20^2+21^2=29^2
400+441=841
841=841